OpenStudy (anonymous):

f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2.

OpenStudy (anonymous):

Generate the first 5 terms of this sequence:

OpenStudy (anonymous):

im sure its easy i just cant figure out whats goin on lol

OpenStudy (btaylor):

so f(3) = f(3-1) + f(3-2) = f(2) + f(1) = 0 + 1 = 1 then f(4) = f(4-1) + f(4-2) = f(3) + f(2) = 1 + 1 = 2 Continue it and you'll get it for f(5). This actually forms a Fibonacci sequence.